OpenStudy (anonymous):

SOMEONE PLEASE HELP!!! Let f(x) =the square root of x, g(x) = x + 8 and h(x) = 9/x. Find the domain of the following function then simplify the expression. (fh)(x)

OpenStudy (jdoe0001):

\(\bf f(x) = \sqrt{x}\qquad g(x)=x+8 \qquad h(x)=\cfrac{9}{x}\\\quad \\ (fh)(x) \implies f(x)\times h(x) \implies \sqrt{x}\times \cfrac{9}{x} \implies \cfrac{9\sqrt{x}}{x}\) the domain will be whatever values for "x" that DO NOT MAKE THE RADICAL to a negative number OR MAKE THE DENOMINATOR to 0

OpenStudy (anonymous):

thank you!!

OpenStudy (jdoe0001):

yw